# Improving Similar Case Retrieval Ranking Performance By Revisiting RankSVM

### v1

@@ Abstract
Given the rapid development of Legal AI, a lot of attention has been paid to one of the most important legal AI tasks–similar case retrieval, especially with language models to use. In our paper, however, we try to improve the ranking performance of current models from the perspective of learning to rank instead of language models. Specifically, we conduct experiments using a pairwise method–RankSVM as the classifier to substitute a fully connected layer, combined with commonly used language models on similar case retrieval datasets LeCaRDv1 and LeCaRDv2. We finally come to the conclusion that RankSVM could generally help improve the retrieval performance on the LeCaRDv1 and LeCaRDv2 datasets compared with original classifiers by optimizing the precise ranking. It could also help mitigate overfitting owing to class imbalance. Our code is available in https://github.com/liuyuqi123study/RankSVM_for_SLR
@@
Information Retrieval, Optimization, Natural Language Processing, Legal Intelligence
@@ 1 Introduction
In recent years, Legal Artificial Intelligence (AI) has attracted attention from both AI researchers and legal professionals([9]). Legal AI mainly means applying artificial intelligence technology to help with legal tasks. Benefiting from the rapid development of AI, especially natural language processing (NLP) techniques, Legal AI has had a lot of achievements in real law applications ([31, 21, 23]).
Among legal AI tasks, similar case retrieval (SCR) is a representative legal AI application, as the appeal to similar sentences for similar cases plays a pivotal role in promoting judicial fairness([21]).
As demonstrated in a lot of work, there are mainly two approaches to do information retrieval. 1) Traditional IR models(e.g.BM25 which is a probabilistic retrieval model([20]) using keywords. 2) More advanced techniques for IR using pre-trained models with deep learning skills, and the latter has achieved promising results in some commonly used benchmark datasets([17]).When using pre-trained language models, there have been a lot of variations trying to incorporate the structural information in legal documents to help with the retrieval([11, 21, 4, 18, 29, 8, 22]) or to utilize LLMs to boost similar case retrieval [32].
However, compared with attention paid to language models, there is less focus on the classifier used for the final ranking. As it is pointed out in an early paper [3], there are two important factors in document retrieval and one of them is that to have high accuracy on top-ranked documents is crucial for an IR system. So we try to look into better classifiers in the hope of improving ranking performance. As a problem of ranking by criterion of relevance, however, SCR is reduced to a 2-class classification problem following a pointwise path in a lot of work, which means the related classifier is only used to produce one label for a query-candidate pair. It is also mentioned by other papers that  [33]the fine-tuned BERT model does not compare which case candidate is more similar to the query case.
In our paper, we reintroduce pairwise methods to do the final ranking. Under pairwise settings, classifiers care about the relative order between two documents, which is closer to the actual ranking task([16]). Among pairwise approaches, we pay extra attention to the Ranking Support Vector Machine(RankSVM) method as a representative method. The advantage of RankSVM is that it aims to minimize the ranking loss and can also mitigate the negative influence of the class-imbalance issue([5, 25]) which is common for SCR tasks. It is also mentioned in some papers ([24]) that the setting of binary labels compared with multi labels could help mitigate the accumulated errors from thresholds learning in RankSVM.
So we are inspired to use RankSVM to explore its performance on similar case retrieval tasks and measure the performance using NDCG as our main metric (Normalized Discounted Cumulative Gain).
@@
In our experiments, we test different kinds of retrieval models when combined with RankSVM to examine their retrieval performance on different similar case retrieval datasets LeCaRDv1([17]) and LeCardv2([15]). Our method is illustrated in Figure 1.
Figure 1: Illustration of Our Method
2 Materials and Methods
2.1 Language Models for SCR
In this paper, we mainly examine three language models: BERT model, BERT+LEVEN, and Lawformer language model, which respectively represent basic language model, knowledge incorporated model, and language model with long inputs. BERT([7]) is a commonly used language model with a maximum window length of 512, and many researchers have done many implementations that have a lot to do with how they understand structural information in legal documents and concatenate them with models that could deal with limited length ([21, 11, 33]). So in our experiments, we both examine the basic BERT model and the BERT model with structural information extracted from legal documents. We use the code from LEVEN([28]) as an example where it incorporates event detection.
Lawformer is a representative model that people propose to deal with longer inputs ([26])which is based on longformer architecture ([2]). Lawformer could process long texts up to respectively 509 for query documents and 3072 for candidate documents.
Some work([27]) moves on to examine another new pre-trained model Mamba([10]) that could deal with long inputs as well, where MAMBA achieves competitive performance compared with transformer-based models using the same training recipe in the document ranking task. So we also measure its performance on the LeCaRDv1 dataset.
2.2 Datasets and Data Preprocessing
@@
In our experiments, we use two benchmark datasets that are commonly used for similar case retrieval tasks–LeCaRDv1 and LeCaRDv2. Here we mainly examine the difference between these two datasets. LeCaRDv1 dataset(A Chinese Legal Case Retrieval Dataset)([17]) and LeCaRDv2 dataset([15]) are two datasets specifically and widely used as benchmarks for similar case retrieval tasks since their release. We first check the difference in the number of query files and candidate files, which is shown in Table 1. While LeCaRDv1 contains 107 queries and 10,700 candidates, LeCaRDv2 contains more queries and candidates. At the same time, it is worth attention that while LeCaRDv1 has a separate candidate pool(folder) of size 100 for each query, there is no subfolder for LeCaRDv2 queries. Moreover, there lies some divergence in their judging criteria. While LeCaRDv1 focuses on the fact part of a case involving key circumstances and key elements, LeCaRDv2 argues that characterization, penalty, and procedure should all be taken into consideration. Even though they say in their paper that there is no explicit mapping function between the Overall Relevance and the sub-relevance, the calculation could be described as follows, see 1.
@@
ratio of relevance candidates 0.1033 0.0004
$relevance_{v2}=relevance_{v1}+relevance_{penalty}+relevance_{procedure.}$ (1)
As it is known, a legal document could usually be partitioned into 3 parts–Parties’ Information, Facts, Holding, and Decision, which is illustrated in Figure 2. Here we follow the common practice to only take fact parts of those documents as inputs, see equation 2 and 3.
Figure 2: Illustration of Structure of A Case
Moreover, we examine the fact parts in both datasets to be used as our inputs, see Figure 3 and Figure 4, from which we could tell that many of them are actually longer than the input limit of the BERT model as well as the Lawformer or MAMBA model. At the same time, as we use the same number of bins in the histogram for both LeCaRDv1 and LeCaRDv2 datasets, it could be seen that while the LeCaRDv2 dataset has query files of longer lengths, it has candidate files of shorter lengths. As it is not
@@
With different benchmark datasets, we conduct different data preprocessing. We also conduct visualization to check LeCaRDv2’s length of documents. See Figure 4 to check their distribution. It could be told that case documents in LeCaRDv2 dataset are shorter .However, as there is no subfolder for LeCaRDv2 dataset in its original repository https://github.com/THUIR/LeCaRDv2, out of consideration for memory and the potential risk of probable overfitting if we conduct experiments with all negative samples among the corpus, we conduct our experiments trying to follow the practice on LeCaRDv1 dataset by building a subfolder of 130 candidates file for each query, among which 30 are those that are labeled relevant which is also similar with the practice taken in LeCaRDv2 baseline where they have negative samples with the ratio of positives and negatives at 1:32. It is shown by our result that it is a wise choice. We also follow different train-test splits. For LeCaRDv1, we follow the 5-fold validation while for LeCaRDv2 we took 640 instances as our training dataset and 160 as our test dataset.
At the same time, the data preprocessing does differ from model to model. For Lawformer, we basically follow the practice in its paper to have candidate files of length 3072 and query files of length 509. For LEVEN code, what we do is we detect their queries and candidate files to get their trigger words and relevant event types. According to the original paper, there are As a result, for non-trigger tokens, it just feeds the sum of token embeddings and position embeddings into the BERT model. For trigger words, we define an event type embedding for each event type and add the corresponding embedding to each input.
$inputx_{without_{e}vent}=func_{tokentoids}(^{\prime}[CLS]^{\prime}+query+^{\prime}[SEP]^{\prime}+cand+^{\prime}[SEP]^{\prime})$ (4)
$inputx_{use_{e}vent}=id_{CLS}+id_{query}+id_{sep}+id_{candidate}+id_{sep}$ (5)
$id_{events}=[0]+id_{query_{e}vent}+[0]+id_{cand_{e}vent}+[0]$ (6)
2.3 Fine-tuning with Pretrained Model And Feature Extraction
When we do fine-tuning, we basically follow the practice taken in LEVEN code https://github.com/thunlp/LEVEN. For BERT-related models, we add a two-class classification layer after the pooled output of BERT denoted as [CLS] and do fine-tuning with the training set.
For Lawformer, the concrete steps are basically the same as BERT. In terms of hyperparameters, we basically follow what is used in LEVEN code since according to our experiments those are better than the original Lawformer code when applied to a smaller batch size. The hyperparameters used are shown in Table 2.
$output=func_{twoclassclassification}([CLS])$ (7)
To extract features, we just use the output of CLS of every query-candidate pair as our features. According to the dimension of CLS, we have 648 features for each query-candidate pair. For further experiments, we only pick the models with the best performance on the validation dataset or test dataset for feature extraction. In the LeCaRDv1 dataset,when training on each fold,we iterate the training dataset for 5 epochs and choose the epoch with the best performance compared with other epochs on the respective validation dataset. After training and validating on each fold, we report the average value for the 5 folds. In the LeCaRDv2 dataset, we iterate the training dataset for 5 epochs and choose the epoch with the best performance on the test dataset to extract features. Our model is only trained from the training data.
The result of RankSVM is used to rank the similarity of candidate-query pairs under the same candidate files.
$features=[CLS]$ (8)
2.4 RankSVM
Since RankSVM is a relatively classic method developed from ordinal-regression SVM to do information retrieval([13, 12]), we try to apply RankSVM([13, 30]) for similar case retrieval task in this work. The mathematical formulation for RankSVM is shown below. Given n training queries ${q_{i}}^{n}_{i=1}$, their associated document pairs $(x_{u}^{(i)},x_{v}^{(i)})$ and the corresponding ground truth label $y_{u,v}^{(i)}$, where a linear scoring function is used without complicated kernel, i.e., $f(x)=w^{T}x$.
The basic form of RankSVM is shown as follows.
$s.t.w^{T}(x_{u}^{(i)}-x_{v}^{(i)})\geq 1-\xi_{u,v}^{(i)},ify_{u,v}^{(i)}=1$ (10)
$\xi_{u,v}^{(i)}\geq 0,i=1,\dots,n.$ (11)
In our experiment, we mainly use the method developed in the paper by Joachims([13]), where the algorithm just has 1 slack variable and it is solved using its dual form. Different algorithms for RankSVM are suitable for different applications. According to our experiment, the result with the algorithm we chose is better than other settings.(e.g., see this work([14]) Also, we use different values of C, and they are 0.001, 0.05, 0.01, 0.02, 0.05, 0.1, 0.5, 1. We report the best result for each model.
It is worth mentioning that here we don’t need a threshold and we just get results in scores and do ranking. Under the setting of predicting with a threshold, as it is said, there will be an implicit presumption that when in training and test, its input has the same data distribution and accumulated error([24]) and may impede model performance.
According to the original LeCaRDv1 and LeCaRDv2 paper, there are 4 classes of relevance. Here we just process it as a two-class classification. When processing labels, as long as two cases are relevant, we view the label of the pair as positive.
@@ 2.4 RankSVM
3 Result
3.1 Results on LeCaRDv1
To test the effectiveness of our method, we conduct our experiments mainly using BERT, BERT+Event, BERT+Event+RankSVM, BERT+RankSVM, Lawformer, Lawformer+RankSVM. The result is shown in Table 3.
From the result on BERT, we could see that RankSVM helps improve performance especially on NDCG, which proves that our model is better at putting highly relevant cases earlier in a similar case list. To better understand the advantage of RankSVM, we finish visualization based on the training and testing dataset on fold 0 to see the score computed by original BERT and BERT+RankSVM. Specifically, we use features extracted by BERT as our input and compress them into two dimensions using t-SNE([19]). After that, we color them according to similarity scores computed by relative models or labels. See Figure 5 and 6 for reference.

### v2

# Woodstock’22: Symposium on the irreproducible science, June 07–11, 2022, Woodstock, NY

@@ Abstract
Given the rapid development of Legal AI, a lot of attention has been paid to one of the most important legal AI tasks – similar case retrieval, especially with language models to use. In our paper, however, we try to improve the ranking performance of current models from the perspective of learning to rank instead of language models. Specifically, we conduct experiments using a pairwise method – RankSVM as the classifier to substitute a fully connected layer, combined with commonly used language models on similar case retrieval datasets LeCaRDv1 and LeCaRDv2. We finally come to the conclusion that RankSVM could generally help improve the retrieval performance on the LeCaRDv1 and LeCaRDv2 datasets compared with original classifiers by optimizing the precise ranking. It could also help mitigate overfitting owing to class imbalance. Our code is available in https://github.com/liuyuqi123study/RankSVM for SLR
@@
[…]
@@ 1 Introduction
In recent years, Legal Artificial Intelligence (AI) has attracted attention from both AI researchers and legal professionals[1]. Legal AI mainly means applying artificial intelligence technology to help with legal tasks. Benefiting from the rapid development of AI, especially within the natural language processing (NLP) domain, Legal AI has had a lot of achievements in real-world law applications[2, 3] such as judgment document generation, legal elements extraction, Legal AI inquiries, and judgment prediction, etc.. Among legal AI tasks, similar case retrieval (SCR) is a representative legal AI application, as the appeal to similar sentences for similar cases plays a pivotal role in promoting judicial fairness and may benefit other applications using large language models(LLM) under settings like RAG[4]. As demonstrated in a lot of work, there are mainly two approaches to do information retrieval. 1) Traditional IR models(e.g.BM25 which is a probabilistic retrieval model[5] based on frequency). 2) More advanced techniques using pre-trained models with deep learning skills. The former method tends to overlook the lexical information within texts as sometimes different words may have the same semantic meaning. The latter, mainly based on transformers[6] can achieve promising results[2], which is also the focus of our work. According to this overview, within the similar case retrieval task or even the legal artificial intelligence, the improvements could be classified into two directions – research based on embeddings and research based on symbols. The first kind focuses on the optimization of word embeddings as many researchers propose that the representation given by pretrained models lacks consideration of legal knowledge within legal texts. The other stresses the extraction of legal relations and events within legal texts, as it helps to improve the interpretability of legal artificial intelligence. When using pre-trained language models, there have been a lot of variations endeavoring in both ways, trying to incorporate the structural information and legal elements in legal documents to help with the retrieval. In BERT_LF, for example, based on representation given by BERT, it also integrates topic information and entity information[7]. In BERT_PLI[8], it designs a model to capture the semantic relationship between paragraphs. In Ma et al.’s work[9], they point out that we need to consider the internal structural information in legal texts especially legal judgments, so they respectively encode the fact part, reasoning part, and sentence part of legal judgments. There is also some work using knowledge graphs instead of language models to encode legal texts, emphasizing the encoding of events to interpretability of the model[10]. Some people have also proposed to utilize LLMs to boost similar case retrieval[11]. In our paper, we instead do not focus on the optimization of language model architecture, but try to optimize classifiers used in the task from a learning-to-rank perspective. Even though the work on the model architecture so far is significant, we want to improve the ranking performance from the problem kind itself. As it is pointed out in early work, there are two important factors to take into account in document retrieval – to rank the most relevant cases more correctly and to consider that different queries may have different numbers of queries[12]. As a problem of ranking by criterion of relevance, however, SCR is reduced to a 2-class classification problem following a pointwise path in a lot of work, which means the related classifier is only used to produce one label for a query-candidate pair without actual comparison. In our work, we are aimed at solving two problems – first to have higher accuracy on top-ranked documents and second to consider each query case by case, in the hope of mitigating class imbalance. In our paper, we reintroduce pairwise methods to do the final ranking. Under pairwise settings, classifiers care about the relative order between two documents, which is closer to the actual ranking task[13]. Among pairwise approaches, we pay extra attention to the Ranking Support Vector Machine(RankSVM) method as a representative method. The advantage of RankSVM is that it aims to minimize the ranking loss and can also mitigate the negative influence of the class-imbalance issue[14, 15] which is common for SCR tasks.
@@
In our experiments, we test different kinds of retrieval models when combined with RankSVM to examine their retrieval performance on different similar case retrieval datasets LeCaRDv1[16] and LeCardv2[17]. Our method is illustrated in Figure1.
Our contribution could be summarized as follows,
  • We first compare different kinds of classifiers and ranking functions used for the final step ranking and give explanations.
  • We first consider SSM architecture in similar case retrieval tasks and give comparisons.
  • We give an efficient practice of how pairwise methods could help improve the ranking performance in SCR task.
Figure 1: Illustraction of our Method
2 Related Work
2.1 Learning-to-rank and Ranking Functions
Learning-to-rank is a long-studied and fruitful field, and people have started thinking about it in similar case retrieval tasks. While it is a common practice to use cross-entropy and fully-connected classifiers for SCR tasks, many researchers have explored different classifiers. For example, it is also mentioned by other papers that[18] the fine-tuned BERT model does not compare which case candidate is more similar to the query case, and they also move on to use a pairwise method borrowing its idea from RankNet[19] on the dataset LeCaRDv1, where the optimization goal is also the number of correctly compared pairs. The difference between RankNet and their method is that they give a different quantity of probability. At the same time, some people also employ InfoNCE in retrieval tasks[20], which is a popular method under a self-supervised setting. Though it considers both positive data and negative data in its optimization goal, it overlooks the relative relationship among different samples. At the same time, as it needs one positive sample during each sampling and always encodes a whole sampling group, it may take a lot of memory and computation compared with pairwise settings. In Ma’s work[9], they also discard the two-class classifier to use LightGBM[21], which originates from gradient boosting decision trees. This is also a pointwise method optimizing the final ranking score. SVM, however, is a relatively old method, but it is still in use in document retrieval. For example, in work by Vacek et al., they integrate SVMs with CNN and GRU in document retrieval [22]. So far, there lacks comprehensive comparison over pairwise method and pointwise method particularly fully-connected classifier for SCR tasks.
2.2 RankSVM
RankSVM, as it is said before, is a representative pairwise method. The framework of the algorithm itself was proposed by Joachims in 1998[23], with the optimization problem it’s aimed at solving built in 2002[24]. It was thought to be able to optimize the retrieval quality with clickthrough data. In search engines, the data could be written in the form of $(q,r,c)$, $q$ is the query, $r$ is the ranking, and $c$ is the clickthrough data finally chosen by the user. For this kind of information retrieval system, it could be formalized as follows. For a query $q$ and a document set $D=\{d_{1},\dots,d_{m}\}$, the best retrieval system would return a ranking $r^{*}$, which corresponds to the relevance between documents in $D$ and query $q$. The evaluation of how good a retrieval function is lies in how its ranking $r_{f(q)}$ can estimate the optimal value. Here $r^{*}$ and $r_{f(q)}$ are both binary relations defined on $D\times D$ and satisfy properties of a weak ordering , which means that $r^{*}\subset D\times D$ and $r_{f(q)}\subset D\times D$ being asymmetric, and negatively transitive. If a document $d_{i}$ is ranked higher than $d_{j}$ for an ordering $r$,i.e. $d_{i}\textless_{r}d_{j}$ then $(d_{i},d_{j})\in r$, otherwise $(d_{i},d_{j})\in r$. In this early paper, it has also talked about the measure of similarity between the system ranking $r_{f(q)}$ and the target ranking $r^{*}$. While average precision is a popular choice, it should be ditched for reasons similar to our view that binary relevance is very coarse. It decides to use Kendall’s $\tau$ as the parameter to be optimized. For $r_{a}\subset D\times D$ and $r_{b}\subset D\times D$, the calculation of Kendall’s $\tau$ is based on the number $P$ of concordant pairs and the number $Q$ of discordant pairs and could be stated as follows
$\tau(r_{a},r_{b})=\frac{P-Q}{P+Q}=1-\frac{2Q}{\binom{m}{2}}$ (1)
Under the setting of binary classification evaluation, maximizing this quantity is minimizing the average ranking of relevant documents. Then, the goal of learning a ranking function is to find a retrieval function $f(q)$, so that the $\tau$ would satisfy
[…]
But a lot of work in machine learning does not consider the form above, instead, they simplify this kind of problem into a binary classification problem. They further explain the problems. First, a ranking function could simply get a really high prediction accuracy by just classifying all data into irrelevant ones as there are always a lot more irrelevant documents. At the same time, sometimes we don’t have an absolute standard of relevance for data. Under many circumstances, users just want to get the most relevant document instead of all relevant documents. Therefore, he proposes an empirical risk minimization approach in his work. Given an independently and identically distributed training sample $S$ of size $n$ containing queries $q$ with their target rankings $r^{*}$
$(q_{1},r_{1}^{*}),(q_{2},r_{2}^{*}),\dots,(q_{n},r_{n}^{*})$ (3)
And the Learner $L$ will select a ranking function f from a family of ranking functions $F$ that maximizes the empirical $\tau$
$\tau_{S}(f)=\frac{1}{n}\sum^{n}_{i=1}\tau(r_{f(q_{i}),r_{i}^{*})}$ (4)
on the training samples. And this optimization goal can be realized by minimizing the number of discordant pairs, and can be approximated by introducing slack variales.
The threshold learning and regularization in RankSVMs have also been well studied. It is also pointed out in some paper [25] that combining binary relevance and Rank SVM has a satisfactory effect as RankSVM could minimize Hamming distance while the binary relevance can remove the need for extra threshold learning. So in our work, we are inspired to use RankSVM with binary relevance.
2.3 Language Models
As we have sketched in the introduction, there has been a lot of work in improving language model architecture. Many researchers have provided their version of BERT[26] variants incorporating legal elements. Lawformer[27] is a meaningful endeavor in pre-trained model breakthroughs. Also, other than the common practice of using transformer-based models, a new model architecture MAMBA[28] has also shown attractive performance, especially for long-input NLP tasks. So in our work, we plan to research four kinds of models: the original BERT model, the BERT model used in the LEVEN baseline dataset, the original Lawformer model, and MAMBA. We wish to use these four kinds of representative models to tell their difference in feature extraction in the SCR task and reach a conclusion about how model architecture may affect the ranking.
3 Methodology
3.1 Overview
In this part, we are to give an introduction to the similar case retrieval task and the steps we take in our model. Similar to most document retrieval tasks, similar case retrieval aims to get candidate cases similar to query cases from a candidate case pool, and they are ranked according to the similarity. Formally, given a query case $q$ and a candidate case set $C=\{c_{i}|i=1,\dots,n\}$, the rank we get for a candidate file $c_{i}\in C$ is $R_{c_{1}/q}$; then the most relevant candidate case is
$c_{*/q}=arg\ max\{R_{c_{i}/q}|c_{i}\in C\}$ (5)
Though there are a lot of parts in a judgment document, we only consider the fact part in a document. Our inputs consist of one query and one candidate. We use language models to extract features from these inputs and send them into our ranking functions.
3.2 Data Preprocessing
In our experiments, the documents we use all come from China Judgements Online; they can usually be partitioned into 3 parts: the information of two parties involved, the facts part, and the decisions made by the court, as is shown in Figure 2. Here we follow the common practice to only take fact parts of those documents as inputs, see equation 6 and 7 respectively for LeCaRDv1 and LeCaRDv2.
Figure 2: Architecture of a legal judgement
3.3 Fine-tuning with Pretrained Model and Feature Extraction
We need language models to capture the deep semantic representation in lagel texts so we conduct fine-tuning first. When we do fine-tuning, we basically follow the framework taken in LEVEN code https://github.com/thunlp/LEVEN. We add a two-class classification layer after the pooled outputs denoted as [CLS] or the hidden state and do fine-tuning with the training set, see equation 9. The hyperparameters used are shown in Table 2.
$output=func_{twoclassclassification}([CLS])$ (8)
As we said before, we decided to use binary relevance labels instead of considering different extents of relevance in our model. According to the description of LeCaRDv1 and LeCaRDv2 papers, relevance could be classified into 4 classes. However, in our experiments, we only consider golden labels( relevance greater than 2) to be positive labels when training, following common practice.
To extract features, we just use the output of CLS of every query-candidate pair as our features. For further experiments, we only pick the models with the best performance on the validation dataset or test dataset for feature extraction. While after training and validating on each fold, we report the average value for the 5 folds on LeCaRDv1, we report the performance for the test dataset on LeCaRDv2. In both datasets,when training on each fold,we iterate the training dataset for 5 epochs and choose the epoch with the best performance compared with other epochs on the respective validation/test dataset.
The features extracted are used to send into ranking functions,stated as follows, see equation 9.
$features=[CLS]$ (9)
3.4 RankSVM
Different forms of RankSVM are suitable for different types of data. In our experiments, we follow the 1-slack algorithm that is solved with the dual problem, which according to our experiments, is better than other settings[23]. The basic form of RankSVM is shown as follows.
$s.t.w^{T}(x_{u}^{(i)}-x_{v}^{(i)})\geq 1-\xi_{u,v}^{(i)},ify_{u,v}^{(i)}=1$ (11)
$\xi_{u,v}^{(i)}\geq 0,i=1,\dots,n.$ (12)
To improve the efficiency, this method also endeavors to choose the direction that could help this iteration find the solution faster. It has also tried to remove those examples that apparently do not affect the support vectors when training.
The efficiency in this algorithm further exists in its cutting-plane solving methods. As all constraints share one slack variant, its bound can easily be verified using its dual problem, written as follows:
$\max_{\alpha\geq 0}\sum_{c\in[0,1]^{n}}\frac{\|c\|_{1}}{n}\alpha_{c}-\frac{1}{2}\sum_{c\in[0,1]^{n}}\sum_{c^{\prime}\in[0,1]^{n}}\alpha_{c}\alpha_{c^{\prime}}x_{c}^{T}x_{c^{\prime}}$ (13)
[…]
It could swiftly get the $\epsilon$-approximate solution, and could be solved within constant times of iteration. The time it takes in classification problem is only $O(sn)$. It is also very efficient for large dataset[29].
To tune how soft the classifier is, we use different values of C, and they are 0.001, 0.05, 0.01, 0.02, 0.05, 0.1, 0.5, 1, 10, 100. In our experiments, we only use the linear kernel. Here we only report the best result for each model to compare with the traditional fully-connected layer.
4 Experiments
4.1 Experiments Settings
we use two benchmark datasets that are commonly used for similar case retrieval tasks–LeCaRDv1 and LeCaRDv2. Here we mainly examine the difference between these two datasets in data distribution. LeCaRDv1 dataset(A Chinese Legal Case Retrieval Dataset)[16] and LeCaRDv2 dataset[17] are two datasets specifically and widely used as benchmarks for similar case retrieval tasks since their release. The details of these two are shown in Table 1. LeCaRDv1 contains 107 queries and 10,700 candidates, among which each query only can retrieve 100 candidates that belong to the respective subfolder. LeCaRDv2 has 800 queries and each query can retrieve from the all-in-one corpus containing 55,192 candidates. At the same time, the ratio of relevant cases is much lower in LeCaRDv2.
@@ 4.1 Experiments Settings
At the same time, it is worth attention that while LeCaRDv1 has a separate candidate pool(folder) of size 100 for each query, there is no subfolder for LeCaRDv2 queries. Moreover, there lies some divergence in their judging criteria. While LeCaRDv1 focuses on the fact part of a case involving key circumstances and key elements, LeCaRDv2 argues that characterization, penalty, and procedure should all be taken into consideration, even though they say in their paper that there is no explicit mapping function between the Overall Relevance and the sub-relevance.
@@
ratio of relevance candidates(positive) 0.1033 0.0004
Moreover, we examine the fact parts in both datasets to be used as our inputs, see Figure 3 and Figure 4, from which we could tell that many of them are actually longer than the input limit of the BERT model as well as the Lawformer or MAMBA model. At the same time, as we use the same number of bins in the histogram for both LeCaRDv1 and LeCaRDv2 datasets, it could be seen that while the LeCaRDv2 dataset has query files of longer lengths, it has candidate files of shorter lengths, which may be due to the way the dataset is built.
@@
With different benchmark datasets, we conduct different data splitting. As there is no subfolder for the LeCaRDv2 dataset in its original repository when our experiments are conducted  [^1] https://github. com/THUIR/LeCaRDv2 and there are only labels for 30 candidate cases concerning a query, we can hardly train our model according to our experiments just using these data. On the other hand, out of consideration for memory and the potential risk of probable overfitting if we conduct experiments with all negative samples among the corpus, we conduct our experiments trying to follow the practice on the LeCaRDv1 dataset by building a subfolder of 130 candidate files for each query, among which 30 are labeled relevant, while the rest are sampled from the large candidate corpus. This is also similar to the practice taken in the LeCaRDv2 baseline experiment where the ratio of positive samples and negative samples is 1:32. It is shown by our result that it is a wise choice. We also follow different train-test splits. For LeCaRDv1, we follow the 5-fold validation, while for LeCaRDv2 we took 640 instances as our training dataset and 160 as our test dataset.
At the same time, the input length does differ among models. For BERT-based models, we truncate the candidate input lengths to 409 and the query input lengths to 100, and the format could be checked in Equation 15 with special tokens. For Lawformer, we basically follow the practice in pretraining to truncate lengths to respectively 3072 and 509, and the format is the same as BERT. For MAMBA, the length of candidate files is 500 and the length of query files is 300,see Equation 16.
$inputx_{BERT}=[CLS]{query}[SEP]{cand}[SEP]$ (15)
$inputx_{Mamba}=document:\{cand\}\backslash n\backslash nquery:\{query\}[EOS]$ (16)
Regarding hyperparameters, we don’t spend a lot of time tuning them and basically follow LEVEN’s code; they could be checked in 2.
@@
4.2 Experiments on LeCaRDv1
To test the effectiveness of our method, we conduct our experiments mainly using BERT, BERT+Event, BERT+Event+RankSVM, BERT+RankSVM, Lawformer, Lawformer+RankSVM,Mamba and Mamba+RankSVM. In our experiments, we take NDCG as our main evaluation metric, which is a common metric in SCR task. It considers the relevance of the result. The result is shown in Table 3.